within eleven years, the age of peter will be half the square of the age he was 13 years ago. calculate the current age of peter...and show your work
Answer by stanbon(75887)

within eleven years, the age of peter will be half the square of the age he was 13 years ago. calculate the current age of peter...
---------------
p-11 = (1/2)(p-13)^2
2p-22 = p^2-26p+169
------
p^2-28p+191 = 0
----
Usable solution to avoid negative age:
p = 16.2361
===============
Cheers,
Stan H.